# How LLMs Aid in UML Modeling: An Exploratory Study with Novice Analysts

### v1

@@ V-A Comparison and Interpretation
Furthermore, by comparing the average scores of the three types of UML models, we can see that LLM’s modeling is more accurate for sequence diagrams. This may be because sequence diagrams involve fewer “relationship” elements and mostly consist of object-type elements that can be directly extracted and created. As a result, LLMs perform quite well in all three criteria for sequence diagrams.
@@ V-B Implication
For educators: The findings of this study inspire a new way to apply LLMs to the education and training of requirements analysis and modeling, providing new directions and goals for future software engineering education. Although at the current stage, using LLMs to generate UML models still has some issues, it can already provide modelers with constructive and critical details and suggestions, which undoubtedly have the potential to be introduced into teaching activities. Using LLMs to evaluate and optimize models created by students is also a promising direction. On the other hand, how to train students to become software engineers who are proficient in using LLMs to assist in modeling work but not overly reliant on it will be a new proposition that software engineering educators need to discuss.
For students: The research results demonstrate that LLMs can already serve as an auxiliary tool to provide effective suggestions, appropriate guidance, and reference examples for software modeling. This can actually be helpful for beginners in software modeling. However, in the process of learning software modeling, students should also be aware of the weaknesses and limitations of LLMs and ensure that they have the ability to correct the occasional errors made by LLMs.
@@ VII Conclusions and Future Work
This paper explored how LLMs assist in creating three types of UML models: use case models, class diagrams, and sequence diagrams, as well as the potential factors affecting their performance. More specifically, we designed a modeling task and asked 45 undergraduate students who participated in a software requirements modeling course to create these three UML models. We required them to use LLM for assistance and submit experimental reports. By collecting and analyzing their reports, we found that LLM can play a role in assisting software modeling but also has shortcomings and limitations. Additionally, we clarified the direction for optimizing LLM’s recognition of modeling elements by comparing the impact of different output formats of UML diagrams.
@@
We hope to expand the sample size to increase the credibility of our conclusions in the future. We also plan to extend the experiments to other types of UML models, such as activity diagrams and state diagrams. Additionally, we aim to conduct controlled experiments comparing the use of LLM for modeling assistance with cases where LLM is not used at all. This quantitative exploration will help us understand the extent to which LLM can assist in improving the quality of UML models.

### v2

@@ V-A Comparison and Interpretation
Furthermore, by comparing the average scores of the three types of UML models, we can see that LLM’s modeling is more accurate for sequence diagrams. This may be because sequence diagrams involve fewer ‘relationship’ elements and mostly consist of object-type elements that can be directly extracted and created. As a result, LLMs perform quite well in all three criteria for sequence diagrams.
@@ V-B Implication
For educators: The findings of this study inspire a new way to apply LLMs to the education and training of requirements analysis and modeling, providing new directions and goals for future software engineering education. Although at the current stage, there are still some challenges of using LLMs to generate UML models, it can give constructive and critical details and suggestions. This is undoubtedly promising to introduce LLMs in teaching activities. That is, using LLMs to evaluate and optimize models manually created by the students is also a promising direction. On the other hand, how to train the students to become software engineers who are proficient in using LLMs to assist in modeling work but not overly reliant on it will be a new proposition that software engineering educators need to discuss.
For students: The results demonstrated that LLMs can serve as an auxiliary tool to provide effective suggestions, appropriate guidance, and reference examples for software modeling. This can actually be helpful for beginners in software modeling. However, in the process of learning software modeling, students should also be aware of the weaknesses and limitations of LLMs and ensure that they have the ability to correct the occasional errors made by LLMs.
@@ VII Conclusions and Future Work
This paper explored how LLMs assist in creating three types of UML models: use case models, class diagrams, and sequence diagrams, as well as the potential factors affecting their performance. More specifically, we designed a modeling task and asked 45 undergraduate students who participated in a software requirements modeling course to create these three UML models. We required them to use LLM for assistance and submit experimental reports. By collecting and analyzing their reports, we found that LLM can play a role in assisting software modeling, but also has shortcomings and limitations. Additionally, we clarified the direction for optimizing LLM’s recognition of modeling elements by comparing the impact of different output formats of UML diagrams.
@@
We hope to expand the sample size to increase the credibility of our conclusions in the future. We also plan to extend the experiments to other types of UML models, such as activity diagrams and state diagrams. Additionally, we aim to conduct controlled experiments comparing the use of LLM for modeling assistance with cases where LLM is not used at all. This quantitative exploration will help us understand the extent to which LLM can assist in improving the quality of UML models. Furthermore, considering the rapid development of large language models in China, such as ERNIE Bot and Qwen, these models will also be included in the scope of future research.
Acknowledgments
This research work has been supported by the National Natural Science Foundation of China (NSFC) with No. 62172311.